Solve using the substitution method:
x-y=5 x+2y=7
I came up with 15y=7

Answer by nyc_function(2741)

x-y = 5...equation 1
x+2y=7....equation 2

x - y = 5
-y = -x + 5

y = (-x + 5)/-1

y = x - 5

Plug (x - 5) for y into equation 2 and solve for x.

x + 2y = 7
x + 2(x - 5) = 7
x + 2x - 10 = 7

3x - 10 = 7

3x = 7 + 10

3x = 17

x = 17/3

We now plug 17/3 into EITHER equation 1 or 2 to find the value of y.

I will use x - y = 5 but you can use the other equation if you'd like.

17/3 - y = 5

-y = 5 - 17/3

-y = -2/3

y = -2/3 divided by -1

y = 2/3
Final answer: x = 17/3 and y = 2/3 which can also be written as a point in the form (17/3, 2/3).
That is the point where the two lines meet on the coordinate plane.
Understand?